# Multi-target stain normalization for histology slides

### v1

@@ Abstract
Index Terms—  Histopathology, stain normalization, domain adaptation, deep learning
@@ 3.1 Experimental setup
We use a UNet++ model with a ResNet18 backbone pretrained on ImageNet. To train the models, we use the Adam optimizer for 150 epochs with a learning rate of 0.001, a batch size of 64, betas of 0.9 and 0.999 respectively, and no weight decay. Our implementation is based on PyTorch [10], and the experiments were run on the Jean-Zay [^1] This work was granted access to the HPC/AI resources of IDRIS under the allocation 2023-AD011013473R1 made by GENCI. cluster on NVIDIA V100 GPUs. The normalization algorithms implementation is based on the torchstain library [11].

### v2

@@ Abstract
Index Terms—  Histopathology, stain normalization, domain adaptation, deep learning.patricio@ubi.pt
@@ 3.1 Experimental setup
We use a UNet++ model with a ResNet18 backbone pretrained on ImageNet. To train the models, we use the Adam optimizer for 150 epochs with a learning rate of 0.001, a batch size of 64, betas of 0.9 and 0.999 respectively, and no weight decay. Our implementation is based on PyTorch [10], and the experiments were run on the Leonardo [^1] We acknowledge the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support. This work was granted access to HPC resources under the allocation IsB28_HyGenAI. cluster on NVIDIA A100 GPUs. The normalization algorithms implementation is based on the torchstain library [11].